# How Contaminated Is Your Benchmark? Quantifying Dataset Leakage in Large Language Models with Kernel Divergence

### v1

@@ Abstract
Dataset contamination, where evaluation datasets overlap with pre-training corpora, inflates performance metrics and undermines the reliability of model evaluations. Quantifying dataset contamination thus becomes essential to ensure that performance evaluations genuinely reflect a model’s ability to generalize to unseen data, rather than relying on memorized examples. To address this problem, we propose Kernel Divergence Score (KDS), a novel method that quantifies dataset contamination by computing the divergence between the kernel similarity matrix of sample embeddings, before and after fine-tuning on the benchmark dataset. Leveraging the insight that fine-tuning affects unseen samples more significantly than seen ones, KDS provides a reliable measure of contamination. Through extensive experiments on controlled contamination scenarios, KDS demonstrates a near-perfect correlation with contamination levels and outperforms existing baselines. Additionally, we perform comprehensive ablation studies to analyze the impact of key design choices, providing deeper insights into the components and effectiveness of KDS. These ablations highlight the importance of leveraging fine-grained kernel-based information and confirm the reliability of the proposed framework across diverse datasets and settings.
@@ 1 Introduction
When a large language model (LLM) performs remarkably well on a benchmark, can you confidently attribute its success to true generalization—or is it simply a reflection of what the model has already seen during pre-training? The reality is, we often don’t know. Beneath the surface of those impressive performance scores lies a critical vulnerability: dataset contamination, a phenomenon where evaluation datasets overlap with the pretraining data of the model (Golchin & Surdeanu, 2024). This overlap artificially inflates reported performance metrics, obscures true generalization capabilities, and raises critical concerns about the reliability of benchmark evaluations. This brings us to a pressing and underexplored question: How can we quantify the degree of dataset contamination?
@@
Addressing this question is crucial to ensuring that performance evaluations genuinely reflect a model’s ability to generalize to unseen data, rather than benefiting from overlap with pretraining data. To formalize the problem, we aim to develop a scoring function $S:(\mathcal{D},\mathcal{M})\rightarrow\mathbb{R},$ that takes a benchmark dataset $\mathcal{D}$ as input and produces a score indicative of its relative contamination level with respect to the given model $\mathcal{M}$. A higher score corresponds to a greater contamination level. Such a score is valuable because researchers can use it to rank multiple benchmarks and prioritize the less contaminated ones, enabling more informed comparisons and reliable evaluation. For the score to be reliable, the scoring function must satisfy two essential properties: monotonicity, which ensures that the score exhibits a positive correlation with the contamination level, and consistency, which means that the score remains stable across independently sampled subsets with the same contamination rate.
@@
Figure 1: Overview of the proposed Kernel Divergence Score (KDS) framework for quantifying dataset contamination in large language models. The process involves extracting sample embeddings from the model before and after fine-tuning on the benchmark dataset $\mathcal{D}$, computing the kernel similarity matrix for each stage, and measuring the divergence between the two matrices $\Phi(Z)$ and $\Phi(Z^{\prime})$. By capturing the changes of embeddings induced by fine-tuning, KDS provides a reliable and interpretable score to quantify the level of dataset contamination.
@@
To quantify dataset contamination, we introduce the Kernel Divergence Score (KDS), which computes the divergence of the kernel similarity matrix of sample embeddings before and after fine-tuning on the benchmark dataset. By analyzing changes in the kernel similarity matrix, KDS captures how fine-tuning reshapes the embeddings for seen and unseen data, providing a more holistic and nuanced perspective on dataset contamination. This approach is motivated by the fact that fine-tuning has a more significant effect on the embeddings of unseen samples, which the model must adapt to, while seen samples exhibit minimal changes due to prior exposure during pre-training. Furthermore, as the proportion of unseen samples increases, their cumulative effect on the kernel divergence score becomes more pronounced. By quantifying these changes, KDS can provide a reliable and interpretable measure of dataset contamination, with scores that proportionally reflect the level of contamination.
@@
  • We propose Kernel Divergence Score, a reliable dataset-level scoring function for quantifying benchmark contamination. To the best of our knowledge, we are the first to leverage the fine-grained information of kernels for scoring contamination levels.
@@ 2 Problem Statement
2.1 Quantifying Benchmark Contamination
The objective is to quantify the relative degree to which a benchmark evaluation dataset, $\mathcal{D}$, has been exposed to the pre-training process of a given LLM, $\mathcal{M}$. In modern LLMs, the pre-training dataset is typically unavailable, making it difficult to directly assess the contamination level. Accordingly, we consider a generalized characterization of the benchmark evaluation data, modeling it as a mixture of both seen and unseen data:
@@ 2.1 Quantifying Benchmark Contamination
which relatively quantifies the contamination of dataset $\mathcal{D}$ with respect to model $\mathcal{M}$. A larger score indicates more contamination and vice versa.
Practical utility.
A reliable scoring function is practically valuable because it allows us to identify benchmark datasets that are less contaminated with respect to model $\mathcal{M}$. By ranking the contamination scores across datasets, we can prioritize benchmark datasets with minimal contamination, ensuring that evaluation results reliably reflect the model’s generalization capabilities rather than memorization of pretraining data. This framework is particularly useful for selecting datasets for fair and trustworthy benchmarking of LLMs. Next, we discuss the desired properties of the scoring function $S$.
@@ 2.2 Reliable Contamination Scores
A comparative study on the contamination level across datasets is reliable only if the scoring function satisfies specific key properties. In this section, we state two essential requirements for a reliable contamination scoring function: Monotonicity and Consistency.
@@
The Monotonicity requirement ensures that the scoring function exhibits a positive correlation with the dataset’s contamination rate, even though the true contamination rate is typically unknown in real-world scenarios. A scoring function satisfying this requirement enables reliable ranking of benchmark datasets for each model based on their contamination scores. The Consistency requirement, on the other hand, ensures that the scores are robust to variations in the specific samples drawn from the same underlying distribution, under the same $\lambda$. This property ensures that the randomness induced from sampling does not substantially affect the overall scoring.
@@ 3 Method: Kernel Divergence Score
Our approach is motivated by the fact that fine-tuning affects the embedding relationships involving unseen samples more significantly than those involving seen samples. For seen samples, the model has already been exposed to similar data during pretraining, leading to minimal shifts in their embedding relationships. In contrast, unseen samples experience more pronounced changes, as the fine-tuning process adjusts the model to better align with the benchmark dataset. By quantifying these changes using the Kernel Divergence Score, we can provide a reliable and granular measure of dataset contamination.
@@
Leveraging the effect of fine-tuning. Our proposed Kernel Divergence Score is based on the kernel matrix before and after fine-tuning. Formally, let $Z^{\prime}\in\mathbb{R}^{n\times d}$ represent the embeddings after supervised fine-tuning on dataset $\mathcal{D}$, where $Z_{i}^{\prime}\in\mathbb{R}^{1\times d}$. Accordingly, we can derive the kernel similarity matrix as:
@@ 5 Ablation Studies
In this section, we conduct an in-depth ablation to understand various design choices of our kernel divergence score.
@@
The results indicate that scoring performance remains consistent across different kernel functions. This suggests that the effectiveness of the Kernel Divergence Score lies in its ability to leverage structural information from kernel representations, rather than being dependent on the specific choice of the kernel. These findings highlight the versatility of our kernel-based approach in quantifying contamination levels across diverse settings.
@@ 6 Discussions
Temporal shift problems of MIA benchmarks.
To ensure the robustness of our approach and mitigate potential biases introduced by temporal shifts, we conducted evaluations using 500 samples from six subsets of the Pile dataset (Gao et al., 2020). The subsets include text data from various sources, including expository proses (Wikipedia), academic papers (PhilPapers), emails (Enron), news articles (HackerNews), web-scraped data (Pile-CC), and user-contributed questions and answers (StackExchange). For each subset, the ‘train’ set is regarded as seen data, while the ‘val’ set serves as unseen data. We mix these two sets according to varying contamination rates to assess our model’s performance under different conditions. This methodology provides a rigorous assessment, ensuring that our model does not exploit temporal cues. As presented in Table 5, the Spearman and Pearson correlation coefficients are both near 1.0, averaging at 0.944 and 0.948, respectively. These findings demonstrate that our method reliably scores contamination levels without relying on temporal shifts.
@@
Our kernel divergence score involves a fine-tuning step to obtain two kernel matrices, followed by the computation of our scoring function. Given a dataset with $N$ samples, the fine-tuning step operates with a time complexity of $O(c_{1}\cdot N)$, where $c_{1}$ is a constant influenced by factors such as average sample length, batch size, and model dimension. The computation of the KDS score has complexity $O(c_{2}\cdot N^{2})$, due to the quadratic nature of kernel matrix operations. In practice, the latency overhead caused by scoring is minimal, as these operations are highly optimized through vectorized computations. As demonstrated in Table 7, the latency measured in seconds for each dataset confirms the efficiency and scalability of our approach. In Appendix E, we confirm that our method can be applied to real-world benchmark datasets, where we employ our method across 11 diverse and widely used benchmarks, most of which have sizes around a few hundred to thousand samples.
@@ 7 Related Works
Data contamination (Magar & Schwartz, 2022; Xu et al., 2024; Balloccu et al., 2024), also known as benchmark leakage, poses a significant challenge in the evaluation of LLMs (Zhou et al., 2023; Duan et al., 2024). To mitigate this problem, one line of research focuses on “decontaminating” datasets by introducing controlled perturbations to reduce overlap with evaluation (Yang et al., 2023). Another line explores methods for detecting contaminated datasets or identifying samples seen during LLM training. Membership inference attack (MIA) techniques (Shokri et al., 2017; Truex et al., 2019) have been employed to classify individual data as seen or unseen (Yeom et al., 2018; Salem et al., 2019; Mattern et al., 2023), with many recent studies specifically targeting LLMs for pre-training data detection (Carlini et al., 2021; Shi et al., 2023; Zhang et al., 2024b; Xie et al., 2024; Li, 2023; Ye et al., 2024). In addition, set-level detection methods have been introduced to identify contamination at a broader dataset level (Oren et al., 2024; Zhang et al., 2024a; Golchin & Surdeanu, 2024). Building on this foundation, our work introduced a novel approach, Kernel Divergence Score, to scoring contamination levels using information derived from embedding kernel similarity matrices. An expanded literature review of MIA is in Appendix F.
@@ Acknowledgement
The authors would like to thank Shawn Im and Seongheon Park for their valuable comments on the manuscript. This work is supported by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation (NSF) Award No. IIS-2237037 & IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, and Philanthropic Fund from SFF.
@@ Appendix D Role of the Normalization Factor
Appendix E A Comprehensive Benchmark Leakage Evaluation In-the-wild
In this section, we provide an extensive evaluation result across large language model and benchmark dataset pairs, based on our Kernel Divergence Score.
Table 10: Evaluation In-the-wild. Kernel Divergence Scores are min-max scaled across datasets. Scores should not be compared across models. Values in parentheses are respective ranks across benchmark datasets.
(Model $\mathcal{M}$, Dataset $\mathcal{D}$) BlQ GSM HSg MTH MPL MPP MHP MNLI OBQ PIQ TQA
Mistral-7B-Instruct-v0.2 0.87 (8) 0.97 (6) 0.81 (9) 1.00 (3) 1.00 (2) 1.00 (4) 0.95 (7) 0.00 (11) 1.00 (1) 0.77 (10) 0.98 (5)
Mistral-7B-Instruct-v0.3 0.89 (7) 0.90 (6) 0.00 (10) 0.98 (4) - (-) 0.93 (5) 1.00 (1) 0.49 (9) 1.00 (2) 0.74 (8) 0.99 (3)
Llama-3.2-1B-Instruct 0.88 (9) 0.97 (5) 0.00 (11) 0.98 (3) 0.92 (6) 0.92 (7) 1.00 (1) 0.58 (10) 0.99 (2) 0.92 (8) 0.98 (4)
Llama-3.2-3B-Instruct 0.35 (10) 0.74 (8) 0.00 (11) 1.00 (1) 0.93 (4) 0.96 (3) 0.98 (2) 0.48 (9) 0.93 (5) 0.77 (7) 0.91 (6)
Llama-3.1-8B-Instruct 0.00 (11) 0.91 (6) 0.55 (8) 1.00 (2) 0.95 (5) 1.00 (1) 0.99 (3) 0.07 (10) 0.90 (7) 0.23 (9) 0.97 (4)
Qwen2.5-1.5B-Instruct 0.56 (9) 0.96 (7) 0.38 (10) 1.00 (1) 0.99 (6) 1.00 (4) 1.00 (2) 0.00 (11) 1.00 (3) 0.83 (8) 0.99 (5)
Qwen2.5-14B-Instruct 0.77 (9) 0.98 (6) 0.24 (10) 1.00 (2) 0.97 (7) 0.98 (5) 0.99 (3) 0.00 (11) 1.00 (1) 0.77 (8) 0.99 (4)
Qwen2.5-7B-Instruct 0.57 (8) 0.99 (7) 0.43 (9) 1.00 (2) 0.99 (6) 1.00 (3) 1.00 (4) 0.28 (10) 1.00 (1) 0.00 (11) 1.00 (5)
Phi-3-small-128k-instruct 0.53 (9) 0.89 (7) 0.10 (10) 1.00 (1) 0.93 (6) 0.97 (4) 0.99 (3) 0.00 (11) 0.99 (2) 0.76 (8) 0.96 (5)
Phi-3-medium-128k-instruct 0.45 (10) 1.00 (6) 0.00 (11) 1.00 (1) 1.00 (4) 1.00 (8) 1.00 (2) 1.00 (2) 1.00 (5) 0.99 (9) 1.00 (7)
Gemma-2-9b-it 0.78 (8) 0.44 (9) 0.35 (10) 1.00 (1) 0.98 (3) 0.93 (6) 0.98 (2) 0.00 (11) 0.96 (5) 0.79 (7) 0.97 (4)
Bloomz-7b1 0.73 (8) 0.96 (5) 0.13 (10) 1.00 (1) 0.95 (7) 0.97 (4) 0.99 (2) 0.00 (11) 0.99 (3) 0.73 (9) 0.96 (6)
Internlm2.5-7b-chat 0.39 (10) 0.89 (6) 0.00 (11) 1.00 (1) 0.93 (5) 0.99 (2) 0.99 (3) 0.56 (9) 0.76 (8) 0.79 (7) 0.95 (4)
Rank Average 8.9 6.5 10.0 1.8 5.1 4.3 2.7 9.6 3.5 8.4 4.8
Score Average 0.60 0.89 0.23 1.00 0.96 0.97 0.99 0.27 0.96 0.70 0.97
Llama Average 0.41 0.87 0.18 0.99 0.94 0.96 0.99 0.38 0.94 0.64 0.95
Mistral Average 0.88 0.94 0.41 0.99 1.00 0.96 0.97 0.24 1.00 0.75 0.99
Qwen Average 0.64 0.98 0.35 1.00 0.98 0.99 1.00 0.09 1.00 0.53 0.99
Phi Average 0.49 0.94 0.05 1.00 0.97 0.98 0.99 0.50 1.00 0.88 0.98
E.1 Benchmark Dataset Details
Here, we list the full names of the benchmark dataset code in Table 10. We also provide dataset details and the specific splits used for evaluation.
  • BlQ is the BoolQ dataset (Clark et al., 2019). It contains binary-choice questions asking about the given passage. We evaluated the validation split that consists of 3270 samples.
  • HSg is the HellaSwag dataset (Zellers et al., 2019). It is a commonsense natural language inference benchmark asking the model to choose an option that best finishes the given sentence. We evaluated the validation split that consists of 10042 samples.
  • OBQ is the OpenBookQA dataset (Mihaylov et al., 2018). It is a 4-way multiple-choice question answering benchmark that requires multi-step reasoning, use of additional common and commonsense knowledge, and rich text comprehension. We evaluated the test split of the ‘main’ subset, which consists of 500 samples.
  • MNLI is the Multi-genre NLI dataset (Williams et al., 2018). It is a crowd-sourced collection of sentence pairs annotated with textual entailment information, covering various spoken and written text. We evaluated the ‘validation_matched’ split that consists of 9815 samples.
  • TFQ is the TruthfulQA dataset (Lin et al., 2022). It consists of question and answer pairs, some of which are correct and others are factually incorrect. Questions are designed in a way that humans may answer falsely due to a false belief or misconception. We evaluated the validation split of the ‘generation’ subset, which consists of 817 samples.
  • PIQ is the Physical Interaction Question-answering dataset (Bisk et al., 2020). It contains questions that require physical commonsense reasoning based on everyday situations. We evaluate the test split of the ‘plain_text’ subset, which consists of 3084 samples.
  • MPP is the Professional Psychology subset of the MMLU dataset (Hendrycks et al., 2021). It contains multiple-choice questions that require expert knowledge in psychology. We evaluate the test split, which consists of 798 samples.
  • MPL is the Professional Law subset of the MMLU dataset (Hendrycks et al., 2021). It contains multiple-choice questions that require expert knowledge in law. We evaluate the test split, which consists of 1101 samples.
  • MHP is the Highschool Psychology subset of the MMLU dataset (Hendrycks et al., 2021). It contains multiple-choice questions that require highschool-level knowledge in psychology. We evaluate the test split, which consists of 544 samples.
  • GSM is the Grade School Math 8K dataset (Cobbe et al., 2021). It contains linguistically diverse grade school math word problems that require multi-step reasoning. Solutions generally involve calculating a series of arithmetic operations. We evaluate the test split of the ‘main’ subset, which consists of 1319 samples.
  • MTH is the MATH-500 dataset (Lightman et al., 2024). It contains a subset of 500 problems from the MATH benchmark, which requires the model to provide a numerical answer to the question. We evaluate the test split, which consists of 500 samples.
E.2 Implementation Details
The implementation details are generally identical to the setup used in our experiments on the WikiMIA, BookMIA, and ArxivTection datasets. We employ 1 epoch of stochastic gradient descent for supervised fine-tuning, with the same LoRA configurations reported in Table 8.
Furthermore, in choosing the target model for evaluation, we tried to choose the latest instruction-tuned version of each model family and size. For instance, the most recent version of 1B and 3B models of the Llama model was 3.2, while it was 3.1 for the 8B model. In such a case, we select the latest version, respectively.
E.3 Results
Evaluation results are provided in Table 10. For better comparison, we min-max scale the scores across benchmark datasets, and their ranks are provided in parentheses. Please note that the scores are not to be compared across models. Overall, the MMLU (Hendrycks et al., 2021), Math500 (Lightman et al., 2024), OBQA (Mihaylov et al., 2018), and TruthfulQA (Lin et al., 2022) datasets showed higher contamination scores, and the two natural language inference datasets, HellaSwag (Zellers et al., 2019) and MNLI (Williams et al., 2018), received lowest scores.

### v2

# How Contaminated Is Your Benchmark? Measuring Dataset Leakage in Large Language Models with Kernel Divergence

@@ Abstract
Dataset contamination, where evaluation datasets overlap with pre-training corpora, inflates performance metrics and undermines the reliability of model evaluations. Measuring dataset contamination thus becomes essential to ensure that performance evaluations genuinely reflect a model’s ability to generalize to unseen data, rather than relying on memorized examples. To address this problem, we propose Kernel Divergence Score (KDS), a novel method that evaluates dataset contamination by computing the divergence between the kernel similarity matrix of sample embeddings, before and after fine-tuning on the benchmark dataset. Leveraging the insight that fine-tuning affects unseen samples more significantly than seen ones, KDS provides a reliable measure of contamination. Through extensive experiments on controlled contamination scenarios, KDS demonstrates a near-perfect correlation with contamination levels and outperforms existing baselines. Additionally, we perform comprehensive ablation studies to analyze the impact of key design choices, providing deeper insights into the components and effectiveness of KDS. These ablations highlight the importance of leveraging fine-grained kernel-based information and confirm the reliability of the proposed framework across diverse datasets and settings. Code is released in https://github.com/deeplearning-wisc/kernel-divergence-score.
@@ 1 Introduction
When a large language model (LLM) performs remarkably well on a benchmark, can we confidently attribute its success to true generalization—or is it simply a reflection of what the model has already seen during pre-training? The reality is, we often don’t know. Beneath the surface of those impressive performance scores lies a critical vulnerability: dataset contamination, a phenomenon where evaluation datasets overlap with the pretraining data of the model (Golchin & Surdeanu, 2024). This overlap artificially inflates reported performance metrics, obscures true generalization capabilities, and raises critical concerns about the reliability of benchmark evaluations. This brings us to a pressing and underexplored question: How can we measure the degree of dataset contamination?
@@
Addressing this question is crucial to ensuring that performance evaluations genuinely reflect a model’s ability to generalize to unseen data, rather than benefiting from overlap with pretraining data. To formalize the problem, we aim to develop a scoring function $S:(\mathcal{D},\mathcal{M})\rightarrow\mathbb{R},$ that takes a benchmark dataset $\mathcal{D}$ as input and produces a score indicative of its relative contamination level with respect to the given model $\mathcal{M}$. A higher score corresponds to a greater contamination level. Such a score is valuable because researchers can use it to rank multiple benchmarks and prioritize the less contaminated ones, enabling more informed comparisons and reliable evaluation. For the score to be reliable, we argue that the scoring function must satisfy two essential properties: monotonicity, which ensures that the score exhibits a positive correlation with the contamination level, and consistency, which means that the score remains stable across independently sampled subsets with the same contamination rate.
@@
Figure 1: Overview of the proposed Kernel Divergence Score (KDS) framework for measuring dataset contamination in large language models. The process involves extracting sample embeddings from the model before and after fine-tuning on the benchmark dataset $\mathcal{D}$, computing the kernel similarity matrix for each stage, and measuring the divergence between the two matrices $\Phi(Z)$ and $\Phi(Z^{\prime})$. By capturing the changes of embeddings induced by fine-tuning, KDS provides a reliable and interpretable score to evaluate the level of dataset contamination.
@@
To measure dataset contamination, we introduce the Kernel Divergence Score (KDS), which computes the divergence of the kernel similarity matrix of sample embeddings before and after fine-tuning on the benchmark dataset. By analyzing changes in the kernel similarity matrix, KDS captures how fine-tuning reshapes the embeddings for seen and unseen data, providing a more holistic and nuanced perspective on dataset contamination. This approach is motivated by the fact that fine-tuning has a more significant effect on the embeddings of unseen samples, which the model must adapt to, while seen samples exhibit minimal changes due to prior exposure during pre-training. Furthermore, as the proportion of unseen samples increases, their cumulative effect on the kernel divergence score becomes more pronounced. By evaluating these changes, KDS can provide a reliable and interpretable measure of dataset contamination, with scores that proportionally reflect the level of contamination.
@@
  • We propose Kernel Divergence Score, a reliable dataset-level scoring function to measure benchmark contamination. To the best of our knowledge, we are the first to leverage the fine-grained information of kernels for scoring contamination levels.
@@ 2 Problem Statement
2.1 Measuring Benchmark Contamination
The objective is to evaluate the relative degree to which a benchmark evaluation dataset, $\mathcal{D}$, has been exposed to the pre-training process of a given LLM, $\mathcal{M}$. In modern LLMs, the pre-training dataset is typically unavailable, making it difficult to directly assess the contamination level. Accordingly, we consider a generalized characterization of the benchmark evaluation data, modeling it as a mixture of both seen and unseen data:
@@ 2.1 Measuring Benchmark Contamination
which measures the relative level of contamination of dataset $\mathcal{D}$ with respect to model $\mathcal{M}$. A larger score indicates more contamination and vice versa.
A reliable scoring function is practically valuable because it allows us to identify datasets that are contaminated with respect to model $\mathcal{M}$. Hence, we now outline the desirable properties that the scoring function $S$ should satisfy.
@@ 2.2 Reliable Contamination Scores
In this section, we state two essential requirements for a reliable contamination scoring function: Monotonicity and Consistency.
@@
The Monotonicity requirement ensures that the scoring function exhibits a positive correlation with the dataset’s contamination rate, even though the true contamination rate is typically unknown in real-world scenarios. A scoring function satisfying this requirement enables reliable ranking of datasets for each model based on their contamination scores. The Consistency requirement, on the other hand, ensures that the scores are robust to variations in the specific samples drawn from the same underlying distribution, under the same $\lambda$. This property ensures that the randomness induced from sampling does not substantially affect the overall scoring.
@@ 3 Method: Kernel Divergence Score
Our approach is motivated by the fact that fine-tuning affects the embedding relationships involving unseen samples more significantly than those involving seen samples. For seen samples, the model has already been exposed to similar data during pretraining, leading to minimal shifts in their embedding relationships. In contrast, unseen samples experience more pronounced changes, as the fine-tuning process adjusts the model to better align with the benchmark dataset. By evaluating these changes using the Kernel Divergence Score, we can provide a reliable and granular measure of dataset contamination.
@@
Leveraging the effect of fine-tuning. Our proposed Kernel Divergence Score is based on the kernel matrix before and after fine-tuning. Formally, let $Z^{\prime}\in\mathbb{R}^{n\times d}$ represent the embeddings after LoRA (Hu et al., 2022) fine-tuning on dataset $\mathcal{D}$, where $Z_{i}^{\prime}\in\mathbb{R}^{1\times d}$. Accordingly, we can derive the kernel similarity matrix as:
@@ 5 Ablation Studies
In this section, we conduct an in-depth analysis to understand various design choices of our kernel divergence score.
@@
The results indicate that scoring performance remains consistent across different kernel functions. This suggests that the effectiveness of the Kernel Divergence Score lies in its ability to leverage structural information from kernel representations, rather than being dependent on the specific choice of the kernel. These findings highlight the versatility of our kernel-based approach in measuring contamination levels across diverse settings.
@@
Table 5: Baseline comparison on PILE data subsets.
Methods Wikipedia PhilPapers Enron HackeerNews Pile_CC StackExchange Average
Zlib 0.861 1.000 1.000 -0.956 -0.782 0.990 0.352
Zlib + FSD 1.000 0.991 0.999 0.323 0.894 0.999 0.868
Perplexity -0.886 0.999 0.999 -0.999 -0.251 0.999 0.144
Perplexity + FSD 1.000 0.990 0.999 0.118 0.908 1.000 0.836
Min-K-0.645 0.996 1.000 -0.955 0.690 0.999 0.348
Min-K+ FSD 0.997 0.952 0.997 0.421 0.908 1.000 0.879
Min-K%++ -0.482 0.960 -0.842 0.561 0.514 0.697 0.235
Min-K%++ + FSD -0.536 0.994 -0.770 0.705 -0.358 0.210 0.041
Kernel Divergence Score (Ours) 0.891 0.982 1.000 0.897 0.895 1.000 0.944
@@ 6 Discussions
Temporal shift problems of MIA benchmarks.
To ensure the robustness of our approach and mitigate potential biases introduced by temporal shifts, we conducted evaluations using 500 samples from six subsets of the Pile dataset (Gao et al., 2020). The subsets include text data from various sources, including expository proses (Wikipedia), academic papers (PhilPapers), emails (Enron), news articles (HackerNews), web-scraped data (Pile-CC), and user-contributed questions and answers (StackExchange). For each subset, the ‘train’ set is regarded as seen data, while the ‘val’ set serves as unseen data. We mix these two sets according to varying contamination rates to assess our model’s performance under different conditions. This methodology provides a rigorous assessment, ensuring that our model does not exploit temporal cues. In Table 5, we provide the baseline comparison results evaluated on the PILE data subsets. On average, our Kernel Divergence Score (KDS) achieves the highest Spearman correlation coefficient. Moreover, as presented in Table 6, both the Spearman and Pearson correlation coefficients are near 1.0, averaging at 0.944 and 0.948, respectively. These findings demonstrate that our method reliably scores contamination levels without relying on temporal shifts.
@@
Our kernel divergence score involves a fine-tuning step to obtain two kernel matrices, followed by the computation of our scoring function. Given a dataset with $N$ samples, the fine-tuning step operates with a time complexity of $O(c_{1}\cdot N)$, where $c_{1}$ is a constant influenced by factors such as average sample length, batch size, and model dimension. The computation of the KDS score has complexity $O(c_{2}\cdot N^{2})$, due to the quadratic nature of kernel matrix operations. In practice, the latency overhead caused by scoring is minimal, as these operations are highly optimized through vectorized computations. As demonstrated in Table 8, the latency measured in seconds for each dataset confirms the efficiency and scalability of our approach.
@@ 7 Related Works
Data contamination (Magar & Schwartz, 2022; Xu et al., 2024; Balloccu et al., 2024; Dekoninck et al., 2024), also known as benchmark leakage, poses a significant challenge in the evaluation of LLMs (Zhou et al., 2023; Duan et al., 2024). To mitigate this problem, one line of research focuses on “decontaminating” datasets by introducing controlled perturbations to reduce overlap with evaluation (Yang et al., 2023). Another line explores methods for detecting contaminated datasets or identifying samples seen during LLM training. Membership inference attack (MIA) techniques (Shokri et al., 2017; Truex et al., 2019) have been employed to classify individual data as seen or unseen (Yeom et al., 2018; Salem et al., 2019; Mattern et al., 2023), with many recent studies specifically targeting LLMs for pre-training data detection (Carlini et al., 2021; Shi et al., 2023; Zhang et al., 2024b; Xie et al., 2024; Li, 2023; Ye et al., 2024). In addition, set-level detection methods have been introduced to identify contamination at a broader dataset level (Oren et al., 2024; Zhang et al., 2024a; Golchin & Surdeanu, 2024). Building on this foundation, our work introduced a novel approach, Kernel Divergence Score, to scoring contamination levels using information derived from embedding kernel similarity matrices. An expanded literature review of MIA is in Appendix F.
@@ 8 Conclusion
Limitations and Future Work
In this work, we introduced the Kernel Divergence Score as a method to reliably measure dataset contamination levels. Compared to sample-wise scoring methods, our kernel divergence score leverages kernels that require quadratic computation. As shown in Table 8, kernel computation itself is not a significant burden in practice. However, it can become a bottleneck in terms of time and space as dataset size increases substantially. To mitigate this computational challenge, we can reduce the amount of computation by adopting local kernel approaches (Segata & Blanzieri, 2010), and by sequentially computing and accumulating the kernel entries to avoid out-of-memory errors. Another promising direction for future research is the application of Positive-Unlabeled (PU) learning (Elkan & Noto, 2008), which focuses on constructing classifiers using only positive and unlabeled data. In the context of dataset contamination detection, PU learning could help in estimating the distribution of contaminated data when only a subset of positive (i.e., contaminated) examples is available. Additionally, exploring kernel calibration methods may enhance the robustness of KDS across different models. By investigating these directions, future work can aim to develop a more universally applicable contamination scoring mechanism, thereby improving the reliability and comparability of contamination assessments across various machine learning models and datasets.
@@ Acknowledgement
The authors would like to thank Shawn Im and Seongheon Park for their valuable comments on the manuscript. Hyeong Kyu Choi, Maxim Khanov, and Yixuan Li are supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, Schmidt Sciences Foundation, and Alfred P. Sloan Fellowship.
@@ Appendix D Role of the Normalization Factor
Appendix E Histogram of Change in Embedding Distances
In Figure 7, we present histograms of the change in embedding distances before and after fine-tuning. We compare three pair types: seen-seen, seen-unseen, and unseen-unseen. The figure reveals that unseen-unseen pairs exhibit more significant shifts, with distances increasing more after fine-tuning overall. This suggests that representations of unseen samples tend to spread farther apart, highlighting and confirmaing our intuition on the distributional shift induced by the fine-tuning process.
Figure 7: Histogram on the values of embedding distance change.